How many Liters of 2.5 M Hydrochloric acid should we use if you have 0.5mol of HCl available.
1
Expert's answer
2020-12-14T14:43:32-0500

Concentration = "\frac{No. of moles of solute }{Volume of solvent}"

Here C is 2.5 M

No.of moles of HCl required is 0.5 mol

Therefore, Volume of solution = "\frac{No. of moles of solute}{Concentration}"

= 0.5/2.5 L = 0.2 L